OpenStudy (anonymous):

12 is 40% of what number? Show work. Giving medel

sammixboo (sammixboo):

\(\rm \color{#20bd23}{\dfrac{part}{whole}=\dfrac{percent}{100}}\) Part = 12 Whole = Unknown Percent = 40 Plug in the values in the formula given above \(\rm \color{#20bd23}{\dfrac{12}{whole}=\dfrac{40}{100}}\) We are solving for the whole, so we can cross multiply 12 and 100, so what is \(\rm \color{#20bd23}{12\times100}\)

OpenStudy (anonymous):

1,200 @sammixboo

sammixboo (sammixboo):

Correct! Now we can take 1,200 and divide it by 40, so what is \(\rm \color{#20bd23}{1,200\div 40}\)

OpenStudy (anonymous):

30 @sammixboo

OpenStudy (anonymous):

oh 30 is the answer right?

sammixboo (sammixboo):

Correct, so we have \(\rm \color{#20bd23}{\dfrac{12}{30}=\dfrac{40}{100}}\) And you can say that 12 is 40% of 30

OpenStudy (anonymous):

ok and 1 more thing @sammixboo

sammixboo (sammixboo):

Okie Dokie :)

OpenStudy (anonymous):

is it ok if i ask the question

sammixboo (sammixboo):

Yeah :)

OpenStudy (anonymous):

How much is a 6% commission of $3,000 in sales